Minimize 0=6X2+2/2, wherex+y (Simplify your answer. Type an exact answer using radicals as needed. Use integers or fra
Answer #1

solution Given datrair Q = 6xtar where at y = 8. =) = 8-k W = ax +2 (8-25 del ferentiate with repect to con both say olem = 1